OpenStudy (anonymous):

Math Question :) Here is the options 3.5 0 3 1?

OpenStudy (anonymous):

What is the y-intercept of the line shown?

OpenStudy (misty1212):

hi

OpenStudy (misty1212):

what do you need, the equation of the line?

OpenStudy (misty1212):

the y intercept is where it crosses the y axis, so i looks to me like it is 3

OpenStudy (anonymous):

I thought so!! THanks!!! Could you help with another? :)

OpenStudy (misty1212):

if you need the entire equation, you need to find the slope right 2, up 1 slope is \(\frac{1}{2}\) making the equation \[y=\frac{1}{2}x+3\] yeah sure if i can i will help

OpenStudy (anonymous):

Okay thanks! Here just a sec

OpenStudy (anonymous):

Two quantities are related, as shown in the table: Which equation best represents the relationship? y = 2x + 11 y = 2x + 10 y = x + 11 y =x + 11

OpenStudy (anonymous):

x y 2 10 4 9 6 8 8 7

OpenStudy (anonymous):

there is the table

OpenStudy (misty1212):

none of the above

OpenStudy (anonymous):

oh okay thanks!

OpenStudy (misty1212):

the rate of change (slope) i\(-1\) since each increase in x has a decrease by 1 in y

OpenStudy (anonymous):

wait then what would it be lol

OpenStudy (misty1212):

i am sorry that is wrong,slope is \(-\frac{1}{2}\)

OpenStudy (misty1212):

should be \[y=-\frac{1}{2}x+11\]

OpenStudy (anonymous):

oh so D?:)

OpenStudy (misty1212):

idk your D did not have -1/2 in it

OpenStudy (anonymous):

oh it does i guess it didnt show up on here lol sorry

OpenStudy (misty1212):

ok then pick that one

OpenStudy (anonymous):

it was right thanks so much!

OpenStudy (misty1212):

hey glad to help